OpenStudy (samigupta8):

The point of intersection of the lines (a^3+3)x +ay+a-3 =0 and (a^5+2)x+(a+2)y+2a+3=0 (a real) lies on the y axis for:- A). No value of a B). Exactly two values of a C). More than 2 values of a D).Exactly one value of a

hartnn (hartnn):

where are you stuck at? equate them, put x=0

OpenStudy (samigupta8):

I found the family of intersection of the two given lines...and then put that x intercept as 0...

hartnn (hartnn):

(a^3+3)x +ay+a-3 =(a^5+2)x+(a+2)y+2a+3 with x =0 gives ay+a = ay +2y +2a+3 isolate y

hartnn (hartnn):

ay+a -3= ay +2y +2a+3**

OpenStudy (samigupta8):

a=-(2)(y+3)

hartnn (hartnn):

isolating y will give y intercept :)

OpenStudy (samigupta8):

A lot ....

OpenStudy (samigupta8):

Except for one value of a

hartnn (hartnn):

y = -(a+6)/2 which value?

OpenStudy (samigupta8):

For a=-6 we will get y intercept 0

hartnn (hartnn):

thats still on y axis

OpenStudy (samigupta8):

Yep....

OpenStudy (samigupta8):

So ans should be for all values of a

hartnn (hartnn):

yes which is same as option C

OpenStudy (samigupta8):

More than 1 value can also be the ans...

OpenStudy (samigupta8):

Bt no option as that...:)

hartnn (hartnn):

A,B, D are definitely false. method of elimination :P